OpenStudy (ashy98):

William raced up a flight of stairs at the rate of 3 stairs per second. The function below shows the total number of stairs f(t) that he is able to climb in t seconds: f(t) = 3t + 7 Which statement best describes the y-intercept of the graph of the function?

OpenStudy (ashy98):

A.) It is equal to 7 and represents the number of stairs William had already climbed before he started racing up B.) It is equal to 3 and represents the number of stairs William can climb in 1 second. C.) It is equal to 4 and represents the number of stairs William had already climbed before he started racing up. D.) It is equal to 3 and and represents the number of stairs William can climb in 1 second.

OpenStudy (academicgurusinc):

The answer is a. The key to this question is to calculate f(t) when t = 0 We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)

OpenStudy (ashy98):

thank youuuu!(:

OpenStudy (ashy98):

could you help me with some more? @AcademicGurusInc

OpenStudy (academicgurusinc):

Sure, what is your question?

OpenStudy (ashy98):

For y = f(x), what is the ordered pair for a point on the graph when x = r - 2? A.) (r - 2, f(r -2)) B.) (x, f(r)) C.) (x, r - 2) D.) (r - 2, f(r))

OpenStudy (academicgurusinc):

This question is testing your understanding of the x and y values of a point within a graph. First step: Understand that an ordered pair is simply (x, y) for a given function. In your case the question states that x = r-2. Therefore, you know that (r-2, y) must hold true. Second step: Solve for y. Since y=f(x) and we know x=r-2, then y = f(r-2). Therefor the ordered pair becomes (r-2), f(r-2). I hope this helps

OpenStudy (academicgurusinc):

Therefore, A is correct

OpenStudy (ashy98):

i need your brain lol Using the function g(x) = 8x - 2, compare and contrast g(-3) and g(3). Which of the following statements is true? A.) The value of g(-3) is larger than the value of g(3). B.) The value of g(-3) is the same as the value of g(3). C.) The value of g(-3) is smaller than the value of g(3). D.) The values of g(-3) and g(3) cannot be compared.

OpenStudy (academicgurusinc):

In this case, the first step requires that you solve for g(-3) and g(3). So let's start with the first one, g(-3): This simply means that x=-3 here so replace the x with -3: g(-3)=8(-3) -2=-24-2=-26 Now solve for g(3) and compare the results to determine your answer. What do find?

OpenStudy (ashy98):

g(-3) =-26?

OpenStudy (ashy98):

so a?

OpenStudy (academicgurusinc):

Not quite, let's try again: your question states that the original function is: g(x) = 8x - 2 so g(-3)= 8(-3) -2= -24-2=-26 Now you do the same thing for g(3) g(3) = 8(3) -2 = 24-2=22 Now which one is greater

OpenStudy (ashy98):

ohhhhh i over looked the (-) g(3) = 8(3) -2 = 24-2=22

OpenStudy (academicgurusinc):

That's right, are you able to determine the answer now?

OpenStudy (ashy98):

c

OpenStudy (academicgurusinc):

Yes, you got it right! Good job! :)

OpenStudy (ashy98):

yay((: okay i have 3 more could you help?(:

OpenStudy (ashy98):

The functions f(x) and g(x) are described below: f(x) = 32x + 8 g(x) = 32x - 9 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units. im 99.9 % sure this is 17

OpenStudy (ashy98):

@AcademicGurusInc

OpenStudy (academicgurusinc):

you are correct!

OpenStudy (ashy98):

sweet((: Which of the following functions best describes a sequence where the first 4 terms are: 7, 4, 1, -2 A.) f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 B) f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 C.) f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 D.)f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1 I "think" its d

OpenStudy (academicgurusinc):

In this case you have to determine the pattern based on the sequence you are given. What pattern do you see?

OpenStudy (ashy98):

1, 7,1, 4, 1 and 1, 7, 1, 3, 1

OpenStudy (academicgurusinc):

Not quite, let's start with the sequence you are given: 7, 4, 1, -2. If you look closely, you will see that by adding -3 to the first number , 7, you will get the second number 4. Similarly, by adding -3 to the second number, 4, you get the third number, 1, and so on. Therefore the pattern requires that you subtract 3 from each number in the sequence to get the next number within the sequence. Does this make sense?

OpenStudy (ashy98):

yes so far

OpenStudy (academicgurusinc):

ok, in that case b must be true. Do you see why?

OpenStudy (ashy98):

by adding -3?

OpenStudy (academicgurusinc):

Yes, "adding -3" is just another way of saying "subtract 3". In other words you are subtracting 3 from every number in the sequence to get the next number. Unfortunately, I will have to leave now. However, I will post the solution to your question on our Tutoring Thursday channel today. Feel free to subscribe to our channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg so that you can be notified once the video has been posted. In the mean time, I hope you found this helpful and good luck!

OpenStudy (ashy98):

thank you so much for your time & help!!!!(:

OpenStudy (academicgurusinc):

Also, feel free to send us your last question and we will add that to the video

OpenStudy (academicgurusinc):

You are quite welcome.

OpenStudy (ashy98):

The table below represents the atmospheric temperature at a location as a function of the altitude: Altitude (in thousand feet) x 15 20 25 30 Temperature (in °C) f (x) 4 -6 -16 -24 The average rate of change of the function between x = 15 to x = 25 is____degrees Celsius per thousand feet and represents the rate of change of temperature per thousand feet. -2?